# InterACT: Inter-dependency Aware Action Chunking with Hierarchical Attention Transformers for Bimanual Manipulation

### v2

@@ 5 Conclusion and Future Work
In this paper, we presented InterACT, a framework for robust bimanual manipulation, which integrates hierarchical attention transformers to capture inter-dependencies between dual-arm joint states and visual inputs. The key contributions of our work include the development of a Hierarchical Attention Encoder and a Multi-arm Decoder. The Hierarchical Attention Encoder aggregates intra-segment information using a segment-wise encoder and integrates inter-segment dependencies through a cross-segment encoder. The Multi-arm Decoder ensures coordinated action sequence generation for each arm through synchronization blocks. Our experimental results, obtained from both simulated and real-world tasks, demonstrate the superior performance of InterACT compared to the baseline ACT framework. The use of CLS tokens, cross-segment encoder and the synchronizatin block significantly enhances the model’s ability to generate accurate and coordinated actions, leading to higher success rates in bimanual manipulation tasks. The ablation studies further highlight the importance of these components in our framework.
While InterACT has shown promising directions in integrating robotic arm joints and visual inputs, it has not yet explored integration with other modalities. Future work will explore integrating additional multi-modal data, such as text or sensory feedback, to further improve the robustness and efficiency of bimanual manipulation tasks. Additionally, addressing the hierarchical nature of tasks will be crucial for better task decomposition and execution. These enhancements could leverage the flexible attention mechanisms demonstrated in this work to manage the added complexity and data integration.
@@
![[Uncaptioned image]](extracted/5856998/images/slot.png)
@@
![[Uncaptioned image]](extracted/5856998/images/plug_insertion.png)
@@
![[Uncaptioned image]](extracted/5856998/images/click_pen.png)
@@
![[Uncaptioned image]](extracted/5856998/images/sweep.png)
@@
![[Uncaptioned image]](extracted/5856998/images/unscrew_cap.png)
@@
Appendix B: Role of CLS tokens
To gain deeper insights into the model’s behavior, we studied how attention weights to CLS tokens change over the timesteps. This study is essential to understand the dynamics of information usage and how the model prioritizes different parts of the input sequence during the decoder phase. Higher attention weights for the CLS tokens indicate a stronger focus on the aggregated information they represent.
The results illustrated in Figure 3 showed that during the phase of interaction between the two arms, significant spikes in attention weights were observed. These spikes occurred at key moments where coordinated actions between the arms were necessary. This indicates that the model heavily relies on the CLS tokens to aggregate and process crucial information when coordinating actions between the arms, highlighting their importance in facilitating precise bimanual manipulation.
Appendix C: Real-robot Setup

### v3

@@ 5 Conclusion and Future Work
In this work, we introduced InterACT, a framework for robust bimanual manipulation, which integrates hierarchical attention transformers to capture inter-dependencies between dual-arm joint states and visual inputs. The key contributions of our work include the development of a Hierarchical Attention Encoder and a Multi-arm Decoder. The Hierarchical Attention Encoder aggregates intra-segment information using a segment-wise encoder and integrates inter-segment dependencies through a cross-segment encoder. The Multi-arm Decoder, while generating action sequence for each arm in parallel, ensures coordinated action sequence generation through synchronization blocks. Our experiment results on both simulated and real-world tasks demonstrate the superior performance of InterACT compared to the baseline ACT. The ablation studies further highlight the importance of these components in our framework. The use of CLS tokens, cross-segment encoder and the synchronization block significantly enhances the model’s ability to generate accurate and coordinated actions, leading to higher success rates in bimanual manipulation tasks.
One of the limitations of InterACT is that the number of CLS tokens and the number of encoder blocks are influenced by the level of coordination required for a given task. While we set a fixed number of CLS tokens across all tasks, finding the optimal hyperparameter requires heavy tuning, as there is no clear rule to guide this selection. Additionally, the level of coordination between the two arms is not straightforward to measure. Coordination can depend on task-specific factors such as temporal synchronization, dependency between arm movements, and precision, which are not easily quantifiable. As a result, determining the optimal hyperparameters becomes even more challenging, requiring manual adjustments and multiple experiments to ensure optimal performance. One potential approach to mitigate this challenge could involve adding an auxiliary loss term that explicitly encourages coordinated actions between the two arms, which might help improve coordination without relying on extensive tuning. However, this may require quantitative definition of coordination and further experimentation. As it stands, the scalability of our approach to more complex tasks or scenarios with diverse coordination requirements remains a challenge.
While InterACT has shown promising directions in integrating robotic arm joints and visual inputs, it has not yet explored integration with other modalities. Future work will explore integrating additional modalities, such as language or sensory feedback, to further improve the robustness of bimanual manipulation tasks. Additionally, addressing the hierarchical nature of tasks will be crucial for better task decomposition and execution. These enhancements could leverage the flexible attention mechanisms demonstrated in this work to manage the added complexity and data integration.
Acknowledgments
We thank the members of the UC Davis Laboratory for AI, Robotics, and Automation (LARA) for their valuable feedback, discussions, and assistance with data collection that contributed to this work.
@@ Acknowledgments
![[Uncaptioned image]](extracted/5930783/figures/slot.png)
@@
![[Uncaptioned image]](extracted/5930783/figures/plug_insertion.png)
@@
![[Uncaptioned image]](extracted/5930783/figures/click_pen.png)
@@
![[Uncaptioned image]](extracted/5930783/figures/sweep.png)
@@
![[Uncaptioned image]](extracted/5930783/figures/unscrew_cap.png)
@@
Appendix B: Real-robot Setup